# On obstructions to the Euler system method for Rankin-Selberg convolutions

### v2

@@ 2 Special elements in the image of the convolution of two modular forms
    There exists $u\in G$ such that $T/(u-1)T$ is free of rank one.
@@
The statements (gI) and (wE) are the two parts of $\mathrm{Hyp}({\mathbb{Q}}(\mu_{p^{\infty}}),T\otimes K)$ in the notation of [14, Section 4.1]. Similarly, (rI) and (sE) constitute in the same way $\mathrm{Hyp}({\mathbb{Q}}(\mu_{p^{\infty}}),T)$. The condition (N) is not mentioned explicitly in [14], but is used in applications of this article to Euler systems, such as [10, Section 11]. Since it is not difficult to study, we include it here.
Let $k\geq 2$. Let $f\in\mathcal{S}_{k}(\Gamma_{1}(N_{f}))$, $g\in\mathcal{S}_{1}(\Gamma_{1}(N_{g}))$ be newforms with characters $\varepsilon_{f}$ and $\varepsilon_{g}$, such that $f$ has no complex multiplication. Let $L,F,H$ be the number fields associated to $f$ in the previous section, and $L_{0}\subset{\mathbb{C}}$ be a number field containing $L$ and all the Fourier coefficients of $g$.
@@
By the work of Deligne and Serre [4, Théorèmes 4.1, 4.6] and [20, Lemma 5.3], there is an odd irreducible Artin representation $\rho_{g}:G_{{\mathbb{Q}}}\rightarrow\operatorname{GL}_{2}(\mathcal{O}_{L_{0}})$ attached to $g$, and $\tilde{\rho}_{g}:G_{{\mathbb{Q}}}\rightarrow\operatorname{PGL}_{2}(L_{0})$ the corresponding projective representation, and we denote $Q=|\tilde{\rho}_{g}(G_{{\mathbb{Q}}})|$. Let $T_{g}$ be a free $\mathcal{O}_{L_{0}}$-module of rank $2$ where $G_{{\mathbb{Q}}}$ acts as $\rho_{g}$. Let $T_{f,g,\mathfrak{p}}=T_{f,\mathfrak{p}}\otimes_{\mathcal{O}_{L_{0}}}T_{g}$. After enlarging $L_{0}$, we assume that $L_{0}$ contains the eigenvalues of every matrix in $\rho_{g}(G_{{\mathbb{Q}}})$.
Remark  The subgroup $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is a finite subgroup of $\operatorname{PGL}_{2}(L_{0})$. Such subgroups admit a well-known classification (see e.g. [23, Proposition 16]) which we now recall, since it is used throughout the paper. More generally, if $K$ is a characteristic zero field, any finite subgroup of $\operatorname{PGL}_{2}(K)$ satisfies exactly one of the following conditions:
@@
Definition  A prime ideal $\mathfrak{p}\subset L_{0}$ is good for $(f,g)$ if the following conditions are satisfied:
@@
    $(T_{f,\mathfrak{p}},\omega_{p}):G_{H}\rightarrow\operatorname{GL}_{2}(\mathcal{O}_{F_{\mathfrak{p}}})\times_{\mathcal{O}_{F_{\mathfrak{p}}}^{\times}}{\mathbb{Z}}_{p}^{\times}$ is surjective, where the fibre product maps are the determinant and the $(k-1)$-th power.
@@
Proof. –  Let $\mathcal{O}$ denote the ring of integers of $F_{\mathfrak{p}}$ and $H(p)$ denote the extension $H(\mu_{p^{\infty}})$ for easier notation. Let $I$ denote the image of $G_{H(p)}$ in $\operatorname{Aut}(T_{f,\mathfrak{p}})\times\operatorname{Aut}(T_{g})$. It is clear that $I\subset\operatorname{SL}_{2}(\mathcal{O})\times\rho_{g}(G_{H})$.
@@
We claim that for any finite Galois extension $M/{\mathbb{Q}}$ unramified at $p$, $G_{{\mathbb{Q}}(\mu_{p^{\infty}})}\rightarrow\mathrm{Gal}(M/{\mathbb{Q}})$ is surjective. Indeed, it is enough to show that ${\mathbb{Q}}(\mu_{p^{\infty}})\cap M={\mathbb{Q}}$. Now, ${\mathbb{Q}}(\mu_{p^{\infty}})\cap M$ is a finite extension of ${\mathbb{Q}}$, so it is equal to ${\mathbb{Q}}(\mu_{p^{n}})\cap M$ for some $n\geq 1$. By assumption, $M/{\mathbb{Q}}$ is unramified at $p$, so ${\mathbb{Q}}(\mu_{p^{n}})\cap M$ is unramified at $p$ over ${\mathbb{Q}}$. Since ${\mathbb{Q}}(\mu_{p^{n}})/{\mathbb{Q}}$ is totally ramified at $p$, so is the intermediate extension $[{\mathbb{Q}}(\mu_{p^{n}})\cap M]/{\mathbb{Q}}$. Thus, ${\mathbb{Q}}(\mu_{p^{\infty}})\cap M$ is both unramified and totally ramified over ${\mathbb{Q}}$ at $p$, so ${\mathbb{Q}}={\mathbb{Q}}(\mu_{p^{\infty}})\cap M$.
Applying this result to the case where $M$ is generated by $H$ and the splitting field of $\rho_{g}$, (which is possible because $p\nmid N_{f}N_{g}$) implies that, for any $\sigma\in G_{{\mathbb{Q}}}$, there is $\sigma^{\prime}\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$ such that $\rho_{g}(\sigma)=\rho_{g}(\sigma^{\prime})$ and $\sigma^{\prime}\in\sigma G_{H}$. In particular, $p_{2}$ is onto and $\rho_{g}(G_{H^{\prime}(\mu_{p^{\infty}})})=\rho_{g}(G_{H^{\prime}})$ for any subextension ${\mathbb{Q}}\subset H^{\prime}\subset H$.
Therefore, it only remains to show that $I$ contains $\{1\}\times\rho_{g}(G_{H})$. To do this, consider the subgroup $G_{1}=p_{2}(\ker{p_{1}})$: it is a normal subgroup of $p_{2}(I)=\rho_{g}(G_{H})$. We can define a continuous surjective homomorphism $\pi:\operatorname{SL}_{2}(\mathcal{O})\rightarrow\rho_{g}(G_{H})/G_{1}$ as follows: for each $x\in\operatorname{SL}_{2}(\mathcal{O})$, pick $y\in\rho_{g}(G_{H})$ such that $(x,y)\in I$, and let $\pi(x)=y\pmod{G_{1}}$. This is well-defined because any two choices of $y$ differ by an element of $G_{1}$. We need to show that $G_{1}=\rho_{g}(G_{H})$.
Recall that, for a Hausdorff topological group $G$, its derived subgroup is the closed subgroup $G^{\prime}$ generated by its commutators, i.e. its elements of the form $ghg^{-1}h^{-1}$ with $g,h\in G$. The subgroup $G^{\prime}$ is preserved by every continuous automorphism of $G$, and it is contained in the kernel of every continuous map $G\rightarrow M$ where $M$ is a Hausdorff topological abelian group.
@@
Thus $\operatorname{SL}_{2}(\mathcal{O})^{\prime}$ contains the normal subgroup generated by the $M_{u}$, whence $\operatorname{SL}_{2}(\mathcal{O})^{\prime}=\operatorname{SL}_{2}(\mathcal{O})$. In particular, every continuous group homomorphism from $\operatorname{SL}_{2}(\mathcal{O})$ to a finite abelian group is trivial.
@@
Proof. –  Let us briefly recall why $\alpha$ exists. For any $\gamma\in\mathrm{Gal}(L/F)$, there is a unique primitive Dirichlet character $\chi_{\gamma}$ such that $(\gamma,\chi_{\gamma})\in\Gamma_{f}$. The map $\gamma\in\mathrm{Gal}(L/F)\longmapsto\chi_{\gamma}(\sigma)\in L^{\times}$ is a cocycle: by Hilbert 90 (for instance [25, §X.1, Proposition 2]), it can be written as $\gamma\longmapsto\gamma(\beta)\beta^{-1}$ for some $\beta\in L^{\times}$. Since $p$ is unramified in $L$, $p$ generates the maximal ideal $\mathfrak{p}\mathcal{O}_{L,\mathfrak{p}}$, so there is some $t\in{\mathbb{Z}}$ such that $\alpha=p^{t}\beta$ lies in $\mathcal{O}_{L,\mathfrak{p}}^{\times}$ and satisfies the required conditions.
By Proposition 2, the image in $\operatorname{Aut}(T_{f,\mathfrak{p}})\times\operatorname{Aut}(T_{g})$ of $\sigma G_{H(\mu_{p^{\infty}})}$ is $I_{\sigma}\times[\sigma\rho_{g}(G_{H})]$, where $I_{\sigma}$ is the image in $\operatorname{Aut}(T_{f,\mathfrak{p}})$ of $\sigma G_{H(\mu_{p^{\infty}})}$. Now, $I_{\sigma}=\begin{pmatrix}\alpha&0\\ 0&\alpha^{-1}\varepsilon_{f}(\sigma)\end{pmatrix}\operatorname{SL}_{2}(\mathcal{O}_{F_{\mathfrak{p}}})$ by [14, Corollary 2.2.3]. $\Box$
@@
The proof makes it clear that the given $\alpha$ works for any element $\sigma^{\prime}\in G_{{\mathbb{Q}}}$ with the same image in $\mathrm{Gal}(H/{\mathbb{Q}})$ as $\sigma$. Furthermore, the element $\alpha$ is defined up to a $\mathfrak{p}$-unit of $F$, i.e. an element of $\mathcal{O}_{F,\mathfrak{p}}^{\times}$.
@@
By [3, Theorem 15.3], $V\otimes W$ is a completely reducible $k[G_{1}]$-module, so the same holds for $U$ by [3, (15.2)], whence by [3, Theorem 15.3] $U$ is a direct sum of irreducible $k[G_{1}]$-submodules. By the Jordan–Hölder theorem [3, (13.7)] applied to the $k[G_{1}]$-module $V\otimes W\simeq V^{\oplus\dim{W}}$, all the irreducible submodules appearing in any composition series of $U$ are isomorphic to $V$, so that $U\simeq V^{\oplus d}$ for some integer $d$: then $d=\frac{\dim_{k}{U}}{\dim_{k}{V}}$. Since $U$ is properly contained in $V\otimes W$, one has $d<\dim_{k}{W}$, so that $d=\dim_{k}\,\mathrm{Hom}_{k[G_{1}]}(U,V)<\dim_{k}{W^{\ast}}$, and $w^{\prime}\in W^{\ast}\mapsto\iota(w^{\prime})_{|U}$ is therefore not injective.
Therefore, $W^{\prime}$ is nonzero, so $W^{\prime}=W^{\ast}$, thus $U=0$ and the conclusion follows.
@@
Proof. –  Proposition 2 shows that $G_{H(\mu_{p^{\infty}})}$ contains an element $\sigma$ acting by $-I_{2}$ on $T_{f,\mathfrak{p}}$ and $I_{2}$ on $T_{g}$, which proves condition (N).
@@
Let $k$ be a finite extension of $\mathcal{O}_{(L_{0})_{\mathfrak{p}}}/\mathfrak{p}$ and assume that $T_{g}/\mathfrak{p}\otimes k$ contains a non-trivial invariant $k[G_{{\mathbb{Q}}}]$-submodule $U$: then $U$ is a line over $k$. Let $c\in G_{{\mathbb{Q}}}$ be the complex conjugation: the line $U$ is stable under $c$, so $U$ is equal to one of the eigenspaces of $c$. Since the characteristic polynomial of $\rho_{g}(c)$ is $(x+1)(x-1)$ and $2\notin\mathfrak{p}$, there is an eigenspace $U_{0}$ for the action of $c$ on $T_{g}/\mathfrak{p}$ such that $U=U_{0}\otimes k$. Then $U_{0}$ is stable under $G_{{\mathbb{Q}}}$, and $T_{g}/\mathfrak{p}$ is reducible under the action of $G_{{\mathbb{Q}}}$.
@@
Suppose that $G_{{\mathbb{Q}}}$ acts reducibly on $T_{g}/\mathfrak{p}$. This action factors through $\rho_{g}(G_{{\mathbb{Q}}})$, and its semisimplification is the direct sum of characters $\alpha,\beta:\rho_{g}(G_{{\mathbb{Q}}})\rightarrow(\mathcal{O}_{L_{0}}/\mathfrak{p})^{\times}$. Thus $\gamma=\alpha/\beta$ is a character of $\rho_{g}(G_{{\mathbb{Q}}})$ and its kernel $K$ consists of matrices $g$ whose reduction modulo $\mathfrak{p}$ is conjugate to $\begin{pmatrix}\alpha(g)&\ast\\ 0&\alpha(g)\end{pmatrix}$.
@@
Hence the image of $K$ in $\operatorname{PGL}_{2}(\mathcal{O}_{L_{0}}/\mathfrak{p})$ is a $p$-group, and $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is an extension of the cyclic subgroup $C_{1}:=\gamma(G_{{\mathbb{Q}}})$ of order prime to $p$ by a $p$-group. Since $p>5$, this is only possible if $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is a non-abelian dihedral group. In this situation, the surjection $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})\rightarrow C_{1}$ factors through the group $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})^{\mathrm{ab}}$ of exponent $2$, so $C_{1}$ has order $2$. Therefore the kernel of $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})\rightarrow C^{\prime}$ is a $p$-group with cardinality $\frac{Q}{2}$, so $\frac{Q}{2}$ is a power of $p$.
@@
Remark  Using the fact that $\operatorname{SL}_{2}(\mathcal{O}_{F_{\mathfrak{p}}})$ is its own derived subgroup, we can slightly refine the proof of Proposition 2 to find $\tau\in G_{{\mathbb{Q}}}^{\prime}$ such that acts on $T_{f,g,\mathfrak{p}}$ by $-\mathrm{id}$. Let now $\eta:G_{{\mathbb{Q}}}\rightarrow\{\pm 1\}$ be a quadratic character and assume that $T_{f,g,\mathfrak{p}}$ satisfies (wE) (resp. (sE)). Let $\sigma\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$ be such that $\ker(\sigma-1\mid T_{f,g,\mathfrak{p}})\otimes(L_{0})_{\mathfrak{p}}$ is a line (resp. $T_{f,g,\mathfrak{p}}/(\sigma-1)T_{f,g,\mathfrak{p}}$ is free of rank one over $\mathcal{O}_{(L_{0})_{\mathfrak{p}}}$).
@@
We are now interested in the properties (wE) and (sE). Suppose that $\sigma\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$ is such that $T_{f,g,\mathfrak{p}}/(\sigma-1)T_{f,g,\mathfrak{p}}$ is free of rank one over $\mathcal{O}_{(L_{0})_{\mathfrak{p}}}$. Let $M\in\operatorname{GL}_{2}(\mathcal{O}_{(L_{0})_{\mathfrak{p}}}),N\in\operatorname{GL}_{2}(\mathcal{O}_{L_{0}})$ be the matrices of the action of $\sigma$ on $T_{f,g,\mathfrak{p}}$ and $T_{g}$ respectively: then $\det{M}$ and $N$ have finite order, and Corollary 2 shows that $M\in\alpha\operatorname{GL}_{2}(\mathcal{O}_{F_{\mathfrak{p}}})$ for a certain $\alpha$ contained in an explicit finite set of $\mathcal{O}_{L,\mathfrak{p}}^{\times}$. It is thus natural to check which matrices $M\in\operatorname{GL}_{2}(\mathcal{O}_{(L_{0})_{\mathfrak{p}}}),N\in\operatorname{GL}_{2}(\mathcal{O}_{L_{0}})$ are such that $N$ and $\det{M}$ have finite order and $\mathcal{O}_{(L_{0})_{\mathfrak{p}}}^{\oplus 4}/(M\otimes N-1)$ is a free $\mathcal{O}_{(L_{0})_{\mathfrak{p}}}$-module of rank one.

### v3

@@ 2 Special elements in the image of the convolution of two modular forms
    There exists $u\in G$ such that T/(u1)TT/(u-1)T is free of rank one.
@@
The statements (gI) and (wE) (resp. (rI) and (sE)) are the two parts of $\mathrm{Hyp}({\mathbb{Q}}(\mu_{p^{\infty}}),T\otimes K)$ (resp. $\mathrm{Hyp}({\mathbb{Q}}(\mu_{p^{\infty}}),T)$) in the notation of [15, Section 4.1]. The condition (N) is not mentioned explicitly in [15], but is used in applications of this article to Euler systems, such as [10, Section 11]. Since it is not difficult to study, we include it here.
Let $k\geq 2$. Let $f\in\mathcal{S}_{k}(\Gamma_{1}(N_{f}))$, $g\in\mathcal{S}_{1}(\Gamma_{1}(N_{g}))$ be newforms with characters $\varepsilon_{f}$ and $\varepsilon_{g}$ such that $f$ has no complex multiplication. Let $L,F,H$ be the number fields associated to $f$ in the previous section and $L_{0}\subset{\mathbb{C}}$ be a number field containing $L$ and all the Fourier coefficients of $g$.
@@
By the work of Deligne and Serre [4, Théorèmes 4.1, 4.6] and [21, Lemma 5.3], there is an odd irreducible Artin representation $\rho_{g}:G_{{\mathbb{Q}}}\rightarrow\operatorname{GL}_{2}(\mathcal{O}_{L_{0}})$ attached to $g$. Let $\tilde{\rho}_{g}:G_{{\mathbb{Q}}}\rightarrow\operatorname{PGL}_{2}(L_{0})$ denote the corresponding projective representation and let $Q=|\tilde{\rho}_{g}(G_{{\mathbb{Q}}})|$. Let $T_{g}$ be a free $\mathcal{O}_{L_{0}}$-module of rank $2$ where $G_{{\mathbb{Q}}}$ acts as $\rho_{g}$. Let $T_{f,g,\mathfrak{p}}=T_{f,\mathfrak{p}}\otimes_{\mathcal{O}_{L_{0}}}T_{g}$. After enlarging $L_{0}$, we assume that $L_{0}$ contains the eigenvalues of every matrix in $\rho_{g}(G_{{\mathbb{Q}}})$.
Remark  The subgroup $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is a finite subgroup of $\operatorname{PGL}_{2}(L_{0})$. Such subgroups admit a well-known classification (see e.g. [24, Proposition 16]) which we now recall, since it is used throughout the paper. More generally, if $K$ is a characteristic zero field, any finite subgroup of $\operatorname{PGL}_{2}(K)$ satisfies exactly one of the following conditions:
@@
Definition  A maximal ideal $\mathfrak{p}\subset\mathcal{O}_{L_{0}}$ is good for $(f,g)$ if the following conditions are satisfied:
@@
    $(T_{f,\mathfrak{p}},\omega_{p}):G_{H}\rightarrow\operatorname{GL}_{2}(\mathcal{O}_{F_{\mathfrak{p}}})\times_{\mathcal{O}_{F_{\mathfrak{p}}}^{\times}}{\mathbb{Z}}_{p}^{\times}$ is surjective, where the fibre product maps are the determinant and the (k1)(k-1)-th power.
Note that, by the results of the previous Section, all but finitely many maximal ideals of $\mathcal{O}_{L_{0}}$ are good.
@@
Proof. –  Let $\mathcal{O}$ denote the ring of integers of $F_{\mathfrak{p}}$. Let $I$ denote the image of $G_{H(\mu_{p^{\infty}})}$ in $\operatorname{Aut}(T_{f,\mathfrak{p}})\times\operatorname{Aut}(T_{g})$. It is clear that $I\subset\operatorname{SL}_{2}(\mathcal{O})\times\rho_{g}(G_{H})$.
@@
We claim that for any finite Galois extension $M/{\mathbb{Q}}$ unramified at $p$, $G_{{\mathbb{Q}}(\mu_{p^{\infty}})}\rightarrow\mathrm{Gal}(M/{\mathbb{Q}})$ is surjective. Indeed, it is enough to show that ${\mathbb{Q}}(\mu_{p^{\infty}})\cap M={\mathbb{Q}}$. Now, ${\mathbb{Q}}(\mu_{p^{\infty}})\cap M$ is a finite extension of $\mathbb{Q}$}, so it is equal to ${\mathbb{Q}}(\mu_{p^{n}})\cap M$ for some $n\geq 1$. By assumption, $M/{\mathbb{Q}}$ is unramified at $p$, so ${\mathbb{Q}}(\mu_{p^{n}})\cap M$ is unramified at $p$ over $\mathbb{Q}$}. Since ${\mathbb{Q}}(\mu_{p^{n}})/{\mathbb{Q}}$ is totally ramified at $p$, so is the intermediate extension $({\mathbb{Q}}(\mu_{p^{n}})\cap M)/{\mathbb{Q}}$. Thus, ${\mathbb{Q}}(\mu_{p^{n}})\cap M$ is both unramified and totally ramified over $\mathbb{Q}$} at $p$, so ${\mathbb{Q}}={\mathbb{Q}}(\mu_{p^{n}})\cap M={\mathbb{Q}}(\mu_{p^{\infty}})\cap M$.
The extension $M/{\mathbb{Q}}$ generated by $H$ and the splitting field of $\rho_{g}$ is unramified at $p$ since $p\nmid N_{f}N_{g}$, so $G_{{\mathbb{Q}}(\mu_{p^{\infty}})}\rightarrow\mathrm{Gal}(M/{\mathbb{Q}})$ is surjective. Hence, for any $\sigma\in G_{{\mathbb{Q}}}$, there is $\sigma^{\prime}\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$ such that $\rho_{g}(\sigma)=\rho_{g}(\sigma^{\prime})$ and $\sigma^{\prime}\in\sigma G_{H}$. In particular, $p_{2}$ is onto and $\rho_{g}(G_{H^{\prime}(\mu_{p^{\infty}})})=\rho_{g}(G_{H^{\prime}})$ for any subextension ${\mathbb{Q}}\subset H^{\prime}\subset H$.
Therefore, it only remains to show that $I$ contains $\{1\}\times\rho_{g}(G_{H})$. The group $p_{2}(I)=\rho_{g}(G_{H})$ contains the normal subgroup $G_{1}=p_{2}(\ker{p_{1}})$. Let $\pi:\operatorname{SL}_{2}(\mathcal{O})\rightarrow\rho_{g}(G_{H})/G_{1}$ be the continuous surjective homomorphism defined as follows: for each $x\in\operatorname{SL}_{2}(\mathcal{O})$, pick $y\in\rho_{g}(G_{H})$ such that $(x,y)\in I$, and let $\pi(x)=y\pmod{G_{1}}$. This is well-defined because any two choices of $y$ differ by an element of $G_{1}$. We need to show that $G_{1}=\rho_{g}(G_{H})$.
Recall that, for a Hausdorff topological group $G$, its derived subgroup is the closed subgroup $G^{\prime}$ generated by its commutators, i.e. its elements of the form $ghg^{-1}h^{-1}$ with $g,h\in G$. The subgroup $G^{\prime}$ is preserved by every continuous automorphism of $G$, and it is contained in the kernel of every continuous map $G\rightarrow A$ where $A$ is a Hausdorff topological abelian group.
@@
Thus $\operatorname{SL}_{2}(\mathcal{O})^{\prime}$ contains the normal subgroup generated by the $M_{u}$, so $\operatorname{SL}_{2}(\mathcal{O})^{\prime}=\operatorname{SL}_{2}(\mathcal{O})$. In particular, every continuous group homomorphism from $\operatorname{SL}_{2}(\mathcal{O})$ to a finite abelian group is trivial.
@@
Proof. –  Let us briefly recall why $\alpha$ exists. For any $\gamma\in\mathrm{Gal}(L/F)$, there is a unique primitive Dirichlet character $\chi_{\gamma}$ such that $(\gamma,\chi_{\gamma})\in\Gamma_{f}$. The map $\gamma\in\mathrm{Gal}(L/F)\longmapsto\chi_{\gamma}(\sigma)\in L^{\times}$ is a cocycle: by Hilbert 90 (for instance [26, §X.1, Proposition 2]), it can be written as $\gamma\longmapsto\gamma(\beta)\beta^{-1}$ for some $\beta\in L^{\times}$. Since $p$ is unramified in $L$, $p$ generates the maximal ideal $\mathfrak{p}\mathcal{O}_{L,\mathfrak{p}}$, so there is some $t\in{\mathbb{Z}}$ such that $\alpha=p^{t}\beta$ lies in $\mathcal{O}_{L,\mathfrak{p}}^{\times}$ and satisfies the required conditions.
By Proposition 2, the image in $\operatorname{Aut}(T_{f,\mathfrak{p}})\times\operatorname{Aut}(T_{g})$ of $\sigma G_{H(\mu_{p^{\infty}})}$ is $I_{\sigma}\times[\sigma\rho_{g}(G_{H})]$, where $I_{\sigma}$ is the image in $\operatorname{Aut}(T_{f,\mathfrak{p}})$ of $\sigma G_{H(\mu_{p^{\infty}})}$. Now, $I_{\sigma}=\begin{pmatrix}\alpha&0\\ 0&\alpha^{-1}\varepsilon_{f}(\sigma)\end{pmatrix}\operatorname{SL}_{2}(\mathcal{O}_{F_{\mathfrak{p}}})$ by [15, Corollary 2.2.3]. $\Box$
@@
It is clear from the proof that $\alpha$ is defined up to a $\mathfrak{p}$-adic unit of $F$ (i.e. an element of $\mathcal{O}_{F,\mathfrak{p}}^{\times}$), and only depends on the image of $\sigma$ in $\mathrm{Gal}(H/{\mathbb{Q}})$.
@@
The irreducible Jordan–Hölder components (cf. e.g. [3, (13.7)]) of the $k[G_{1}]$-module $V\otimes W$ are isomorphic to $V$, so the same holds for $U$. Now, $V\otimes W\simeq V^{\oplus\dim{W}}$ is a completely reducible $k[G_{1}]$-module, so $U$ is a direct sum of irreducible $k[G_{1}]$-submodules by [3, (15.2), Theorem 15.3]. Therefore $U\simeq V^{\oplus d}$ as a $k[G_{1}]$-module, with $d=\frac{\dim_{k}{U}}{\dim_{k}{V}}$. Since $U$ is properly contained in $V\otimes W$, one has $d<\dim_{k}{W}$, so that by Schur’s lemma $d=\dim_{k}\,\mathrm{Hom}_{k[G_{1}]}(U,V)<\dim_{k}{W^{\ast}}$. Hence $w^{\prime}\in W^{\ast}\mapsto\iota(w^{\prime})_{|U}$ is not injective, thus $W^{\prime}$ is nonzero. Therefore $W^{\prime}=W^{\ast}$, thus $U=0$ and the conclusion follows.
@@
Proof. –  By Proposition 2, $G_{H(\mu_{p^{\infty}})}$ contains an element $\sigma$ acting by $I_{2}$ on $T_{f,\mathfrak{p}}$ and $I_{2}$ on $T_{g}$, which proves condition (N).
@@
Let $k$ be a finite extension of $\mathcal{O}_{(L_{0})_{\mathfrak{p}}}/\mathfrak{p}$ and assume that $T_{g}/\mathfrak{p}\otimes k$ contains a non-trivial invariant $k[G_{{\mathbb{Q}}}]$-submodule $U$: then $U$ is a line over $k$. Let $c\in G_{{\mathbb{Q}}}$ be the complex conjugation: the line $U$ is stable under $c$, so $U$ is equal to one of the eigenspaces of $c$. Since the characteristic polynomial of $\rho_{g}(c)$ is (x+1)(x1)(x+1)(x-1) and $2\notin\mathfrak{p}$, there is an eigenspace $U_{0}$ for the action of $c$ on $T_{g}/\mathfrak{p}$ such that $U=U_{0}\otimes k$. Then $U_{0}$ is stable under $G_{{\mathbb{Q}}}$, and $T_{g}/\mathfrak{p}$ is reducible under the action of $G_{{\mathbb{Q}}}$.
@@
Suppose that $G_{{\mathbb{Q}}}$ acts reducibly on $T_{g}/\mathfrak{p}$. This action factors through $\rho_{g}(G_{{\mathbb{Q}}})$, and its semisimplification is the direct sum of characters $\alpha,\beta:\rho_{g}(G_{{\mathbb{Q}}})\rightarrow(\mathcal{O}_{L_{0}}/\mathfrak{p})^{\times}$. Thus $\gamma=\alpha/\beta$ is a character of $\rho_{g}(G_{{\mathbb{Q}}})$ and its kernel $K$ consists of matrices $g$ whose reduction modulo $\mathfrak{p}$ is conjugate to $\begin{pmatrix}\alpha(g)&\ast\\ 0&\alpha(g)\end{pmatrix}$. Note that $\gamma$ sends the complex conjugation to 1-1, so its image has even order.
@@
Hence the image of $K$ in $\operatorname{PGL}_{2}(\mathcal{O}_{L_{0}}/\mathfrak{p})$ is a $p$-group, and $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is an extension of the cyclic subgroup $C_{1}:=\gamma(G_{{\mathbb{Q}}})$ of order prime to $p$ by a $p$-group. Since $p>5$, this is only possible if $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is a non-abelian dihedral group. In this situation, the surjection $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})\rightarrow C_{1}$ factors through the group $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})^{\mathrm{ab}}$ of exponent $2$, so $C_{1}$ has order $2$. Therefore the kernel of $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})\rightarrow C_{1}$ is a $p$-group with cardinality $\frac{Q}{2}$, so $\frac{Q}{2}$ is a power of $p$.
@@
Remark  Using the fact that $\operatorname{SL}_{2}(\mathcal{O}_{F_{\mathfrak{p}}})$ is its own derived subgroup, we can slightly refine the proof of Proposition 2 to find $\tau\in G_{{\mathbb{Q}}}^{\prime}$ such that acts on $T_{f,g,\mathfrak{p}}$ by $\mathrm{id}$. Let now $\eta:G_{{\mathbb{Q}}}\rightarrow\{\pm 1\}$ be a quadratic character and assume that $T_{f,g,\mathfrak{p}}$ satisfies (wE) (resp. (sE)). Let $\sigma\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$ be such that $\ker(\sigma-1\mid T_{f,g,\mathfrak{p}})\otimes(L_{0})_{\mathfrak{p}}$ is a line (resp. $T_{f,g,\mathfrak{p}}/(\sigma-1)T_{f,g,\mathfrak{p}}$ is free of rank one over $\mathcal{O}_{(L_{0})_{\mathfrak{p}}}$).
@@
We are now interested in the properties (wE) and (sE). Suppose that $\sigma\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$ is such that $T_{f,g,\mathfrak{p}}/(\sigma-1)T_{f,g,\mathfrak{p}}$ is free of rank one over $\mathcal{O}_{(L_{0})_{\mathfrak{p}}}$. Let $M\in\operatorname{GL}_{2}(\mathcal{O}_{(L_{0})_{\mathfrak{p}}}),N\in\operatorname{GL}_{2}(\mathcal{O}_{L_{0}})$ be the matrices of the actions of $\sigma$ on $T_{f,\mathfrak{p}}$ and $T_{g}$ respectively. Then $\det{M}$ and $N$ have finite order, and Corollary 2 shows that $M\in\alpha\operatorname{GL}_{2}(\mathcal{O}_{F_{\mathfrak{p}}})$ for a certain $\alpha$ contained in an explicit finite set of $\mathcal{O}_{L,\mathfrak{p}}^{\times}$. It is thus natural to check which matrices $M\in\operatorname{GL}_{2}(\mathcal{O}_{(L_{0})_{\mathfrak{p}}}),N\in\operatorname{GL}_{2}(\mathcal{O}_{L_{0}})$ are such that $N$ and $\det{M}$ have finite order and $\mathcal{O}_{(L_{0})_{\mathfrak{p}}}^{\oplus 4}/(M\otimes N-1)$ is a free $\mathcal{O}_{(L_{0})_{\mathfrak{p}}}$-module of rank one.